# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 3.9 Aalto Datasets (2018; 2019)
The Aalto University desktop [35] and mobile [90] datasets are large-scale datasets collected using an online typing test on desktop computers and mobile devices. The desktop dataset has 136 million keystrokes gathered over three months from 168,000 subjects. Participants were instructed to transcribe fifteen (15) sentences in english by typing them as fast and accurate as possiable. The 15 sentences were randomly taken fom a set of 1,525 samples consisting of at least 3 words and a maximum of 70 characters. Subjects were allowed to make typing errors, correct them or add new characters when typing, and as a result, they could type more than 70 characters. The dataset is categorized as a controlled free-text dataset because it involves transcribing - subjects did not type contents of their own but were shown what to type. There were 15 sessions during each participants transcribed only one sentence. The front end of the online test was deployed as an HTML web page leveraging CSS for website styling. JavaScript was used for dynamic events on the website such as keypresses and the data were stored in a MySQL database. A key-logger was used in capturing keystroke data including imtestamps of key up and down events as well as the key that was pressed.
The Aalto mobile dataset [90] was collected from 37,370 participants in a preforming transcription services on a website. It is an extension of the Aalto desktop dataset and followed the same procedure but on mobile devices. Subjects were to memorize a given sentence, and then type it quickly and accurately. Being a web-based method with browser-side logger, as opposed to laboratory-based, there was less control over the quality of the data collected, and web applications on mobile devices have limited access privileges. As a result, many participants’ data were found with undefined keycodes (229 or 0); many devices’ down and up keystroke events are were equal, producing invalid keystroke duration of $<$10ms; and the keycodes of some pressed keys are missing.
@@ 3.10 AR Dataset (2021)
Wahab et al. [125] used a new dataset of 500,000 keystrokes, designed for account recovery, to determine the performance of five differing keystroke dynamics-based algorithms [52], [60], [66], [65] - Euclidean distance, Manhattan distance, Scaled Manhattan distance, Mahalanobis distance, and the Gunetti and Picardi’s algorithm. A total of 44 subjects were used during data collection including both students and university employees. The subjects had to complete an account recovery form that contained different fields. This type of data is called semi-fixed as it is neither free-text nor fixed-text. Two collection sessions were hosted for gathering the data. In the first session, the subjects had to fill out an enrollment form ten times. The data collected from session one was used to create the profile of a subject. In the second session, the subjects had to fill out the same form again five times. The data collected from session two was used as genuine keystroke data for the corresponding subject and imposter data for other subjects. There was a one or two weeks gap between the sessions. The fields of the enrollment form were- Full name, Address, City, Zip, Phone, Email, Declaration, Password. For each of the 42 subjects the enrollment task was completed ten times. 28 subjects attended the second session and 16 subjects acted as imposter. Five algorithms were used for the evaluation of this account recovery dataset. Scaled Manhattan distance algorithm showed the best performance. The best EER was 5.47% for individual fields, 0% for five fields combined and 0% for seven fields combined.
@@ 3.11 Multi-K dataset (2022)
Wahab et al. [124] created two novel free-text keystroke dynamics datasets. The first was a multi-keyboard keystroke dataset. Data was collected by leveraging four physical keyboards - mechanical, membrane, ergonomic and laptop keyboards. For the second dataset, data was collected in both English and Chinese languages. 86 users participated in the data collection - 60 users for the multi-keyboard dataset and 26 users for the bilingual dataset. A keylogger deployed on a website was leveraged for the collection of both datasets.
To examine the impact of multi-keyboard and multi-language data in keystroke dynamics, Wahab et al. employed two free-text algorithms: the Instance-based Tail Area Density (ITAD) metric and the D-Vectors model. On average, each user provided around 14,000 keystrokes for the multi-keyboard dataset. For each user, per keyboard type, data of the initial visit was used for enrollment, while data from the second visit was used for testing. There were 200 keystrokes in each test sample, and the performance was measured using the Equal Error Rate (EER). The results indicated that enrolling and testing with different keyboards affect keystroke dynamics performance where keyboard size and layout were two significant factors.
@@ 3.12 Touch Screen and Mobile Datasets
El-Abed, Dafer and El Khayat [37] coined a new benchmark named RHU Keystroke which uses touch screen keystorke dynamics. They used a Windows Phone app for data collection. The data were stored in a keystroke dynamics database and necessary features were extracted. The participants in the data collection were students of varying backgrounds to ensure the scenario was as real as possible. Their goal was to make a publicly available benchmark for keystroke authentication. A total of 53 participants participated in 3 sessions typing the password rhu.university 15 times in each session. On average there was 5 days gap between two sessions. The RHU Keystroke benchmark consisted of four timing features: PP, PR, RP, RR.
@@
Stragapede et al. [111] presented a new database of mobile Human-Computer Interaction (HCI) named BehavePassDB. 81 users participated in the data collectin sessions through a dedicated mobile app installed on the user’s device. The user was asked to perform a series of tasks. In order to check whether the model can differentiate between the notion of user and device, two dedicated sessions were carried out by another person (imposter) on the same exact device at the time of data collection. Data were collected from 15 background sensors as well, i.e., touchscreen, accelerometer etc. For the evaluation purpose, Long-Short Term Memory (LSTM) architecture was used.
@@
In [8], Al-Obaidi and Al-Jarrah proposed a classifier leveraging a median-based method for keystroke authentication on mobile devices. The classifier was used as an anomaly detector on the SU dataset [15]. For this dataset, the CMU password (".tie5Roanl") was used. There were both timing and touch features of mobile devices in the dataset. There were 11 characters in the password including the Enter key resulting in 71 features in total (timing and touch screen features). 42 users took part in the data collection typing the same password 51 times. 34 entries were used for training and 17 were used for testing. The data collection was performed on Android devices, a tablet and a mobile phone. The proposed statistical classifier model used the median of the training features. The testing features were authenticated against the training features. There were three timing features: Hold (H), Up-Down (UD), Down-Down (DD) and two touch screen features: Pressure (P) and Finger Area (FA). The first five entries of each subject were used to create an imposter set. Al-Obaidi and Al-Jarrah achieved an EER of 8.53% for 41 timing features and an EER of 6.79% for both timing and touch screen.
@@
In [7], Al-Obaidi and Al-Jarrah proposed a keystroke authentication system for use on mobile devices. They adopted an anomaly detector leveraging statistical distance-to-median that used both the timing and touch screen features. The goal was to develop a suitable keytroke dataset for mobile devices and evaluate the performance of the distance-to-median anomaly detector. Al-Obaidi and Al-Jarrah performed their experiment on Nexus smartphones and tablets and used five features for the authentication system - Hold, Latency or Up-Down (UD), Down-Down (DD), Pressure (P), Finger Area (FA). The system consisted of two phases. In the training module, the required tasks were: 1) Registration of users through user-id and password, 2) Selection of the number of attempts of password entry and the pass-mark, 3) Password typing for the selected number of times, 4) Template generation. In the testing phase, the test vector was compared against the template stored in the database. The test vector was generated from the login attempt. The data was collected from 56 users each typing a 10 character password 51 times. For a different pass-mark per subject, they achieved an EER of 0.049. For a global pass-mark, Al-Obaidi and Al-Jarrah achieved an EER of 0.054 and implemented three verification models based on distance to obtain their results. Al-Obaidi and Al-Jarrah obtained an FRR (false rejection rate) of 5.6% for the FAR (false acceptance rate) of 5%.
@@
Kim and Kang [67] proposed a novel keystroke dynamics-based authentication method for mobile devices named FACT (free text, accelerator, coordinate, and time). In order to evaluate their proposed model, Kim and Kang collected a novel keystroke dataset based on mobile devices. They used an Android-based application for the data collection. Total 50 subjects participated in the data collection. 20 free-text scripts were typed by each subject where each script contained about 200 keystrokes. Special characters were not included in the scripts. The subjects provided the data in both English and Korean. 17 features were extracted from the raw data including not only the typing pattern but also hand size, finger length, fingertip size, and muscle flexibility. The proposed FACT method achieved an EER lower than 1%. Kim and Kang made their collected keystroke dataset public [67].
@@ 4 Keystroke Authentication Algorithms
To guage the effectiveness of keystroke authentication systems, an authentication algorithm is required. Different kinds of keystroke authentication algorithms have been developed by researchers. In this section state-of-the-art keystroke authentication algorithms are explored.
4.1 Statistical Algorithms
Statistical algorithms are easy to implement, model and explain. They are deterministic and also do not require heavy computing power and/or memory. However, they require manual extraction of features which could be difficult and/or daunting and are less accurate when dealing with very large datasets. Below, we describe common statistical algorithms used for keystroke dynamics.
4.1.1 Gaussian Mixture Model
Gaussian Mixture Model (GMM), a linear non-Gaussian multivariate statistical method, is a popular algorithm used for handling non-Gaussian data. It is a statistical method based on the weighted sum of probability density functions of multiple Gaussian distributions [122, 34]. GMM generates a vector of mean values corresponding to each component and a matrix of covariance which includes components’ variances and the co-variances between each other. The parameter set $\lambda$ is used for expressing GMM. The parameter $\lambda$ comprises of component weights $w\textsubscript{$i$}$, mean vector $\overrightarrow{\mu\textsubscript{$i$}}$ and covariance matrix $\Sigma\textsubscript{$i$}$ as shown in equation $\lambda=\{w_{i},\overrightarrow{\mu_{i}},\Sigma_{i}\},\text{$i=1,...,M$}$. The parameters are estimated using the iterative expectation–maximization (EM) algorithm [33]. In every iteration, parameter ($\lambda$) is updated if the iteration yields a higher likelihood and fits the distribution of the training dataset.
Given a set of reference samples $\lambda$ (also known as profile), the mixture density for the keystroke input sample $\overrightarrow{x}$ is defined as the weighted linear combination of M pure Gaussian distributions as shown in equation - $p(\overrightarrow{x}|\lambda)=\sum_{i=1}^{M}w_{i},p_{i}(\overrightarrow{x})$; where $p_{i}(\overrightarrow{x})=\frac{1}{(2\pi)^{D/2}|\Sigma_{i}|^{1/2}}exp\biggl{\{}-\frac{1}{2}(\overrightarrow{x}-\overrightarrow{\mu})^{T}\Sigma_{i}^{-1}(\overrightarrow{x}-\overrightarrow{\mu})\biggl{\}}$
4.1.2 Euclidean Distance
[…]
4.1.3 Manhattan and Scaled Manhattan Distance
[…]
(a) Sample below the median
(b) Sample above the median
Figure 5: Graphical representation of how the ITAD metric is computed from the PDF and CDF. In (a), the sample is below the median and the ITAD metric is determined by the left tail area of the PDF. If the left tail area is x, then x point of the CDF (on the left) indicates the ITAD metric score. In (b), the sample is above the median and the ITAD metric is determined by the right tail area of the PDF. If the right tail area is y, then (1-y) point of the CDF (on the right) indicates the ITAD metric score.
4.1.4 Instance-Based Tail Area Density
[…]
The illustration in Figure 5 depicts the computation of the ITAD metric using a graphical representation. The ITAD metric is calculated as the area of the tail of the PDF. If the sample falls below the median, the ITAD metric is determined by the tail area on the left, while if it is above the median, it’s determined by the tail area on the right. ITAD performs better than any distance metrics (such as the Scaled Manhattan) which are based on mean and standard deviation of graphs and are affected by outliers. ITAD metric scores ranges between 0 and 0.5 and can be interpreted as a measure of similarity. The larger the score is, the closer the test sample is to the profile.
Similarity $\text{Similarity Score}=\frac{1}{N}\sum_{i=1}^{N}S_{\text{$i$}}$ (6)
4.1.5 Mahalanobis Distance
Mahalanobis distance is an example of a distribution-based algorithm. It is an effective multivariate distance metric that measures the distance between a point and a distribution [96, 132, 131, 66]. It takes into consideration the correlation between features which is more suitable when features are not independent. It is best for finding multivariate outliers and finds its applications in multivariate anomaly detection, classification and one-class classification. The Mahalanobis equation is - $D=(X-\mu_{t})\cdot C^{-1}\cdot(X-\mu_{t})$; where $X$ is the feature vector of a test sample, $\mu_{t}$ is the mean feature calculated from the profile template, and $C^{-1}$ is the inverse covariance matrix of the user profile template.
4.1.6 Gunetti and Picardi’s Metric
Gunetti and Picardi’s free-text algorithm [52] combines typing speed (A-measure) and the degree of disorder (R-measure) to measure similarity. The ‘A’ measure represents the distance between typing samples $S_{1}$ and $S_{2}$ in terms of $n$-graphs, where $n$ is the number of consecutive keystrokes ($n=1$ is monograph, $n$=2 is digraph, $n$=3 is trigraph etc.), is given in equation -
[…]
The ‘R’ measure on the other hand quantifies the degree of disorder between two sequences. Given a array $V$ of $K$ elements, a simple measure of the degree of disorder of $V$ with respect to $V^{\prime}$ (its ordered counterpart) is computed as the sum of the distances between the position of each element in $V$ and the position of the same element in $V^{\prime}$.
4.2 Machine Learning-Based Algorithms
Machine learning methods have been increasingly employed for keystroke dynamics. Here, we will detail the state-of-the-art machine learning algorithms with a focus on keystroke dynamics.
4.2.1 Supervised Learning
[…]
Supervised Machine Learning can be further divided into classification and regression. Classification uses an algorithm to accurately predict a class label on a test data - e.g., classifying user behavior as genuine or imposter. Binary classification refers to a problem involving two classes and multi-class classification refers to a problem with greater than three classes. On the other hand, regression is used for predicting the relationship between the label and input variables. Here, instead of predicting classes, regression makes numerical predictions such as predicting the weather or average salary. In keystroke dynamics, the focus is either to verify if the user is who he/she claimed to be (known as authentication), or to identify the user who the query sample belongs to (known as identification). Both scenarios belong to the classification type, hence, keystroke dynamics uses the classification techniques.
4.2.1.1 Support Vector Machine
A support vector machine (SVM) is a type of supervised learning model leveraged for classification problems. The SVM algoritum computes hyperplane(s) that uniquely seperate classes of points of data. The hyperplane, also known as the decision boundary, is an N-dimensional space plane, where N is the number of features in the dataset. Of several hyperplanes possible, the goal is to compute a boundry with the maximum margin that separates the classes (legitimate users from impostors) [46, 69, 128, 107]. This algorithm derives its name from support vectors, which are data points that are closer to the hyperplane and decides the orientation of the plane.
In keystroke dynamics, it is assumed that the keystroke samples in the instance of enrollment and verification have the an equal distribution, which is not usually the case, and as a result a lower prediction accuracy is obtained from most traditional machine learning classifiers. In this regard, SVM has a support for adaptive learning [26], where an existing SVM classifier (source) is adapted to another SVM classifier (target) based on some projection techniques. Knowing that there exist variations in users’ typing pattern at every sample, especially when the samples have a wide time margin (weeks, months or years apart), the adaptive SVM modifies the hyperplane of a pre-trained classifier using the new data points with minimum retraining, and therefore performs better than the regular SVM trained on the same amount of labelled data. The SVM classifier is a commonly used classifier in keystroke dynamics, especially on a small feature set dataset [24, 44].
4.2.1.2 Naive Bayes
The Naive Bayes classifier is a probabilistic machine learning model used for classification and based on the Bayes theorem with “naive" independent assumptions between features [16, 56, 31]. The Bayes’ theorem finds the probability of an event $a$ occurring given that the probability of another event $B$ has already occurred, as shown in equation $P(a|B)=\frac{P(B|a)\times P(a)}{P(B)}$; where $P(a|B)$ is the posterior probability, $P(a)$ is the prior probability of the label class, $P(B)$ is the prior probability of the predictor and $P(B|a)$ is the likelihood. The Naive Bayes classifier offers several advantages, such as being less computationally demanding compared to other traditional machine learning models, and performing well on small training data when the assumption of feature independence holds. Additionally, it can be used for solving multi-class prediction problems, which are identification problems that involve predicting one of several possible classes (or outcomes) for a given input. In keystroke dynamics, where dependencies between features may exist, the Naive Bayes classifier may not be as effective and is therefore less commonly used.
4.2.1.3 Tree-Based Models
Tree-based models are a type of supervised machine learning algorithm used for classification. They use a set of conditional statements to recursively split training data into subsets. The end result model is a roadmap of logical decisions that describes the dataset. Tree-based models are easy to implement and interpret, less computationally intensive, and are popularly used in keystroke dynamics because of their non-parametric attributes, making it preferable even for a non-Gaussian distributed data [104]. Tree-based models are tolerant to outliers and intra-class variation problems, work with both numerical and categorical variables, and require less data preprocessing steps such as variable transformations (scaling, normalization etc). However, they are prone to overfitting and high variance. Popular tree-baseed methods used in keystroke dynamics are random forest [17, 11], gradient boost and extreme gradient boost (XGBoost) [107, 27]
4.2.1.4 K-Nearest Neighbors
K-Nearest Neighbors (KNN) is a non-parametric supervised machine learning classifier (although can also be unsupervised) that classifies new data points based on their proximity with data points of already known classes. KNN works on the assumption that data points of similar classes are closer to each other and can be found near each other [58, 109, 16]. At testing, During testing the distances between a new point of data and prior points (neighbors) is calcualted. A class label is assigned to a new data point based on a majority vote between its K nearest neighbors, where K is the number of neighbors that will be checked to determine the class to assign for the new query point. Generally, the optimal value of K is found experimentally, as a low value of K results in low bias but high variance, while a high value of K gives high bias and low variance. Distance metrics (such as the Euclidean distance, Manhattan distance, Hamming distance etc.) are used in calculating the distance between the query point and other data points. These distance metrics form the decision boundaries. KNN is known as a lazy learner as it does the nearest neighbor calculation at every run time. Because of this, it is used only on less dimensional data and has high computational cost when used on high dimensional data.
4.2.1.5 Multi-Layer Perceptron
Multi-layer Perceptron (MLP) is a type of supervised learning algorithm cosisting of input, output, and several hidden layers (those in between the input and output layer). The input data to be processed are passed to input layer, the output layer does the classification, and the hidden layer (non-linear layer) preforms the the computational work. Therefore, data flows in forward direction from the input layer to the output layer. MLP is a neural network with multiple layers and each layer is a combination of neurons such that the output of some neurons are passed as input to other neurons in the next layer. The number of neurons and hidden layers used is dependent on the type of classification problem being solved. MLP is capable of learning non-linear models, but it requires several hyperparameter tuning such as the number of hidden neurons, hidden layers and iterations. Hence, it is computationally intensive and requires both genuine and imposter data to train. Neural network is capable of producing much better performance than the statistical algorithms and any other machine learning algorithms when working with big data that has both genuine and impostor samples [5, 53, 113, 92].
4.2.2 Unsupervised Learning
Unsupervised learning is a type of ML used for identifying patterns in datasets that contain unlabeled data points whereas with supervised learning the data is labeled and the algorithm is trained on this data to make predictions. In keystroke dynamics, unsupervised learning algorithms are often used for outlier detection. Some of the unsupervised learning algorithms used in keystroke dynamics include One-class SVM (OCC), K-NN and K-Means [62, 125]. These algorithms are useful in situations where obtaining labeled data is difficult, costly, or impossible, and the ground truth (label) is unknown.
@@ 4.2.2 Unsupervised Learning
Table 3: A summary of keystroke authentication algorithms by the names and references of the algorithms, the references of the datasets that were used to evaluate the algorithms and the results obtained using these algorithms.

### v3

@@ 3.9 Aalto Datasets (2018; 2019)
The Aalto University desktop [35] and mobile [90] datasets are large-scale datasets collected using an online typing test on desktop computers and mobile devices. The desktop dataset has 136 million keystrokes gathered over three months from 168,000 subjects. Participants were instructed to transcribe fifteen sentences in English by typing them as fast and accurately as possible. The 15 sentences were randomly taken from a set of 1,525 samples consisting of at least 3 words and a maximum of 70 characters. Subjects were allowed to make typing errors, correct them, or add new characters when typing, and as a result, they could type more than 70 characters. The dataset is categorized as a controlled free-text dataset because it involves transcribing, subjects did not type contents of their own but were shown what to type. There were 15 sessions and participants transcribed only one sentence during each session. The front end of the online test was deployed as an HTML web page leveraging CSS for website styling. JavaScript was used for dynamic events on the website such as keypresses and the data were stored in a MySQL database.
The Aalto mobile dataset [90] was collected from 37,370 participants performing transcription services on a website. It is an extension of the Aalto desktop dataset and follows the same procedure but on mobile devices. Subjects were to memorize a given sentence, and then type it quickly and accurately. Being a web-based method with a browser-side logger, as opposed to laboratory-based, there was less control over the quality of the data collected, and web applications on mobile devices have limited access privileges. As a result, many participants’ data were found with undefined keycodes; many devices’ down and up keystroke events were equal, producing invalid keystroke duration of $<$10ms; and the keycodes of some pressed keys were missing.
@@ 3.10 AR Dataset (2021)
Wahab et al. [125] collected a new dataset of 500,000 keystrokes, designed for account recovery, to determine the performance of five differing keystroke dynamics-based algorithms [52], [60], [66], [65]: Euclidean distance, Manhattan distance, Scaled Manhattan distance, Mahalanobis distance, and the Gunetti and Picardi’s algorithm. A total of 44 subjects were used during data collection including both students and university employees. The subjects had to complete an account recovery form that contained different fields. This type of data is called semi-fixed as it is neither free-text nor fixed-text. Two collection sessions were hosted to gather the data. In the first session, the subjects had to fill out an enrollment form ten times. The data collected from session one was used to create the profile of a subject. In the second session, the subjects had to fill out the same form again five times. The data collected from session two was used as genuine keystroke data for the corresponding subject and imposter data for other subjects. There was a one or two-week gap between the sessions. The fields of the enrollment form were- Full name, Address, City, Zip, Phone, Email, Declaration, Password. 28 subjects attended the second session and 16 subjects acted as imposter. Five algorithms were used for the evaluation of this account recovery dataset. The Scaled Manhattan distance algorithm showed the best performance. The best EER was 5.47% for individual fields, 0% for five fields combined, and 0% for seven fields combined.
@@ 3.11 Multi-K dataset (2022)
Wahab et al. [124] created two novel free-text keystroke dynamics datasets. The first was a multi-keyboard keystroke dataset. Data was collected by leveraging four physical keyboards: mechanical, membrane, ergonomic, and laptop keyboards. For the second dataset, data was collected in both English and Chinese languages. 86 users participated in the data collection: 60 users for the multi-keyboard dataset and 26 users for the bilingual dataset. A keylogger deployed on a website was leveraged for the collection of both datasets.
To examine the impact of multi-keyboard and multi-language data in keystroke dynamics, Wahab et al. employed two free-text algorithms: the Instance-based Tail Area Density (ITAD) metric and the D-Vectors model. On average, each user provided around 14,000 keystrokes for the multi-keyboard dataset. For each user, per keyboard type, data from the initial visit was used for enrollment, while data from the second visit was used for testing. There were 200 keystrokes in each test sample, and the performance was measured using the Equal Error Rate (EER). The results indicated that enrolling and testing with different keyboards affect keystroke dynamics performance where keyboard size and layout were two significant factors.
@@ 3.12 Touch Screen and Mobile Datasets
El-Abed, Dafer, and El Khayat [37] developed a new benchmark named RHU Keystroke for touch screen keystroke dynamics. They used a Windows Phone app for data collection. The data was stored in a keystroke dynamics database and necessary features were extracted. The participants in the data collection were students of varying backgrounds to ensure the scenario was as real as possible. A total of 53 participants participated in 3 sessions typing the password rhu.university 15 times in each session. On average there was a 5 days gap between successive sessions. The RHU Keystroke benchmark consists of four timing features: PP, PR, RP, and RR.
@@
Stragapede et al. [111] created a new database of mobile Human-Computer Interaction (HCI) named BehavePassDB. 81 users participated in the data collection sessions through a dedicated mobile app installed on the user’s device. The user was asked to perform a series of tasks. In order to check whether the model can differentiate between the notion of user and device, two dedicated sessions were carried out by another person (imposter) on the same exact device at the time of data collection. Data was collected from 15 background sensors as well, i.e., touchscreen, accelerometer, etc. For evaluation, Long-Short Term Memory (LSTM) architecture was used.
@@
Al-Obaidi and Al-Jarrah [8] proposed a classifier leveraging a median-based method for keystroke authentication on mobile devices. The classifier was used as an anomaly detector on the SU dataset [15]. For the SU dataset, the CMU password (“.tie5Roanl”) was used. There were both timing and touch features of mobile devices in this dataset. There were 11 characters in the password including the Enter key resulting in 71 features in total (timing and touch screen features). 42 users took part in the data collection typing the same password 51 times. 34 entries were used for training and 17 for testing. Data collection was performed on Android devices, a tablet, and a mobile phone. The proposed statistical classifier model used the median of the training features. The testing features were authenticated against the training features. There were three timing features: Hold (H), Up-Down (UD), Down-Down (DD), and two touch screen features: Pressure (P) and Finger Area (FA). The first five entries of each subject were used to create an imposter set. Al-Obaidi and Al-Jarrah achieved an EER of 8.53% for 41 timing features and an EER of 6.79% for both timing and touch screen features.
@@
Al-Obaidi and Al-Jarrah [7] proposed a keystroke authentication system for use on mobile devices. They adopted an anomaly detector leveraging statistical distance-to-median that used both the timing and touch screen features. The goal was to develop a suitable keystroke dataset for mobile devices and evaluate the performance of the distance-to-median anomaly detector. They performed their experiment on Nexus smartphones and tablets and used five features for the authentication system: Hold, Latency or Up-Down (UD), Down-Down (DD), Pressure (P), and Finger Area (FA). The system consisted of two phases. In the training module, the required tasks were: 1) Registration of users through user-id and password, 2) Selection of the number of attempts of password entry and the pass-mark, 3) Password typing for the selected number of times, and 4) Template generation. In the testing phase, the test vector was compared against the template stored in the database. The test vector was generated from the login attempt. The data was collected from 56 users each typing a 10-character password 51 times. For a different pass mark per subject, they achieved an EER of 0.049. For a global pass-mark, Al-Obaidi and Al-Jarrah achieved an EER of 0.054 and implemented three verification models based on distance to obtain their results. They obtained an FRR (false rejection rate) of 5.6% for the FAR (false acceptance rate) of 5%.
@@
Kim and Kang [67] proposed a novel keystroke dynamics-based authentication method for mobile devices named FACT (free text, accelerator, coordinate, and time). In order to evaluate their proposed model, Kim and Kang collected a novel keystroke dataset based on mobile devices. They used an Android-based application for the data collection. A total of 50 subjects participated in the data collection. 20 free-text scripts were typed by each subject and each script contained about 200 keystrokes. Special characters were not included in the scripts. The subjects provided the data in both English and Korean. 17 features were extracted from the raw data including not only the typing pattern but also hand size, finger length, fingertip size, and muscle flexibility. The proposed FACT method achieved an EER lower than 1%. Kim and Kang made their collected keystroke dataset public [67].
@@ 4 Keystroke Authentication Algorithms
The authentication algorithm is a crucial part of any keystroke-based authentication system. This section describes state-of-the-art keystroke authentication algorithms. A summary of keystroke authentication algorithms is shown in Table 3.
@@
Table 3: A summary of keystroke authentication algorithms, the datasets that were used to evaluate the algorithms, and the results obtained using these algorithms.